Both you and your older brother would like to have $26,500 in 10 in years. Because of your success in this class, you feel that you are a more savvy investor than your brother and will be able to earn an annual return of 11.5 percent compared to your brother's 9.8 percent. How much less than your brother will you have to deposit today
Answer #1

Amount required to be deposited = Future Value/(1+Interest Rate)^Number of years

Amount required to be deposited by me = 26,500/(1.115)^10 = $8,922.72

Amount required to be deposited by brother = 26,500/(1.098)^10 = $10,404.53

Hence, less amount i.e. difference = 10,404.53 - 8,922.72

= $1,481.81